Solve the following system of equations for the unknown variables.
x + 2y = 14
x – 3y = 9

Answer by mamiya(56)

x + 2y = 14
x – 3y = 9
there are many ways to approach this, i am going to use the substitution method.
for that method, you have to take one equation , write letter in term of the other letter and plug that in the other equation.
let take the first one, x+2y =14 --> x = 14 - 2y
by plugging that in the second equation, we get
x- 3y = 9 --> 14-2y -3y =9
--> 14-5y = 9
--> -14 +14 -5y = -14 +9
--> -5y= -5
--> -5y/(-5) = -5/(-5)
--> y=1
and we know that x = 14 -2y = 14 - 2(1)= 14-2 = 12
so the answer to the system is ( 12, 1)